### v1

@@ 3.3 Summary of Collected Data
Fig. 2(a) demonstrates the influence of vehicle type on yielding rates. The results indicate that SUVs and sedans are less likely to yield compared to other vehicle types, whereas buses and trucks exhibit the highest yielding rates. We classified pedestrians into eight distinct categories to examine differences in driver interactions, as illustrated in Fig. 2(b): A: pedestrian, B: person on a vehicle such as a scooter, C: mix of pedestrian types, D: other, E: person with a dog, and F: person with a stroller or child. The most common type is Category A and the second most common one is pedestrian with a dog (Category E). As exhibited by Fig. 2(c), the drivers are more likely to yield to a group of pedestrians. In this study, interactions between cyclists and vehicles have been excluded, though the interactions are kept in the published dataset for completion. Fig. 2(d) illustrates the variations in yielding behavior (yielding rate) at different hours of the day. Our findings indicate that drivers are more likely to yield during noon compared to other times throughout the day.
@@
We collected variables that related to the site location to reflect location characteristics so that the results can be used to improve pedestrian walkability. For instance, Fig. 3(a) demonstrates that if there is at least one bus stop present, the number of bus stops correlates with yielding rates. We discovered that the driver-yielding rate is higher when more bus stops are near the intersection. Unintuitively, this trend does not hold true for intersections without any bus stops in close proximity. Regarding the number of lanes on the major road (Fig. 3(b)), drivers are less likely to yield when there are more lanes. Fig. 3(c) reveals a correlation between lower driver-yielding rates and the presence of trees obstructing the drivers’ view of waiting pedestrians. Specifically, the results indicate that the highest driver-yielding rates occur at intersections without any trees blocking the crosswalk, while the lowest yielding rate is found at intersections where tree cover is present at most of the intersection corners.
@@
Markings play a crucial role in increasing driver yielding, as demonstrated in Fig. 4(a). Intersections with markings exhibit a higher percentage of driver yielding compared to driver non-yielding. Interestingly, sites with standard markings (two parallel lines) had higher yielding rates than those with continental markings, contrary to our expectations. Among the sites, 12 had no markings, 3 had standard markings, and 3 had continental markings. It is important to note the small sample size of sites with standard and continental markings. The presence of a bike lane correlated with a decrease in driver yielding, although only 9.1% of the interactions in the dataset occurred at sites with bike lanes, as shown in Fig. 4(b). As expected, sites with crosswalk signs had higher driver yielding rates, as depicted in Fig. 4(c).
(a) Type of markings vs. yielding.
(b) Presence of bike lanes vs. yielding.
(c) Presence of crosswalk sign vs. yielding.
@@
Lastly, various adjacent land use types were documented during data extraction, such as distance to a park or school and the presence of multi-family housing. Fig. 6 illustrates the impact of two of these land uses—restaurants/bars and parking lots. Sites featuring restaurants/bars and/or parking lots constituted a significant portion of the dataset, and these land uses were associated with a higher rate of driver yielding. However, in the case of parking lots, there are still more non-yielding events.
(a) Presence of parking lots vs. yielding.
(b) Presence of restaurants/bars vs. yielding.
@@
3.4 Summary of Statistics
Table 3 presents the summary statistics for pedestrian behavior at the 18 intersections studied. Factors include the number of total interactions, the number of close calls, and the driver’s yielding rate. Intersection 16 had the most interactions with 840 events, and intersection 11 had the fewest number of interactions with 20. While intersection 16 also had the highest yielding rate of 70.36%, intersection 11 had the lowest yielding rate of 0. The total number of pedestrians also evidences this at each location since the driver-yielding rates are often higher at these locations with more pedestrians. Another noteworthy observation is that the maximum waiting time for pedestrians at certain intersections approaches 1 minute. In some instances, pedestrians were forced to halt midway through crossing, potentially elevating the risk of crashes. The diminished yielding rate at unsignalized, unmarked intersections with low pedestrian volume may be attributed to drivers’ lack of anticipation for pedestrian presence.
@@ 3.4 Summary of Statistics
Sites Total interactions Close calls Ped Stop Midway Yielding rate (%) Total num of Ped. Average time (intend to cross - start to cross) (s) Average time (start to cross - finish cross) (s) Max waiting time (s) Min waiting time (s)
@@ 4 Modeling Driver Yielding
This section outlines the statistical analysis conducted to model driver-yielding rates at 18 unsignalized intersections under study. By utilizing the developed model, we can identify significant factors influencing driver yielding at these sites and estimate the likelihood of a driver yielding under the prevailing conditions at similar locations.
@@ 4.1 Logistic Regression
Logistic regression is a robust technique for binary classification that yields a model with good interpretability [44]. It models the probability of a binary dependent variable $Y$, which in this case is the driver-yielding response that has a binary outcome - either the driver yields to the pedestrian or fails to yield. We assume a linear relationship between the independent variable and the log-odds of the event. This linear relationship is shown in (1), where $L$ is the log-odds, $\beta_{i}$ are the model parameters, and $P(Y=1)$ for driver yielding and $1-P(Y=1)$ for driver not yielding are the probabilities of the aforementioned binary outcomes. The logistic regression model takes the form of Equation (2) to estimate the probability of the driver yielding $P(Y=1)$, and conversely, Equation (3) estimates the probability of the driver not yielding $1-P(Y=1)$. For the driver-pedestrian interactions considered, the outcome is either driver-yielding ($Y=1$) or non-yielding ($Y=0$).
@@
where $\beta_{0}$ is the constant, $\beta_{i}$ is the coefficient of the explanatory variable, $x_{n}$ is the predictor variable, and $n$ is the number of features.
The collected interaction data were divided into training and testing data sets, with 80% of the data used for training the model and the remaining 20% reserved as hold-out test data. The parameters used in the logistic model were selected through the stepwise logistic method, which is a technique used to identify the critical explanatory variables. The stepwise process combines forward and backward elimination to select the crucial variables [45]. The selection criterion for these predictor variables relies on the Akaike Information Criterion (AIC) [46], a measure that gauges both prediction error and model quality. Recognizing that some variables in this dataset are categorical, we implemented feature engineering to convert these into dummy variables, adopting values of either 0 or 1.
@@ 4.2 Model for Driver-pedestrian Interaction
Table 4.2 lists the variables identified as significant and thus included in the logistic regression model of driver yielding. Based on the variable selection criterion introduced above, the vehicle speed, whether opposite direction yield, the width of major crossing road, the presence of restaurants/bars/parking lots, the distance to the nearest park, and the distance to the nearest school are all significant. Therefore, the logistic regression model presented in Table 4.2 is constructed using the above variables. The coefficients $\beta_{i}$ within the model are determined using maximum-likelihood estimation, an iterative procedure aimed at minimizing the error in the predicted probabilities. These coefficients $\beta_{i}$ articulate the magnitude of the influence each predictor variable has on the binary outcome of driver yielding or not yielding. Additionally, the standard error provides an estimate of each coefficient’s variability, while the Z-score conveys the deviation from the mean in units of standard deviation. A test statistic, the p-value, reflects the probability of obtaining extreme results from the statistical analysis. Specifically, a low p-value signifies a predictor’s likely relevance to the model.
@@
Variables coefficient std err z P>|>|> |z|normal-||| Effect
@@
Based on the result in Table 4.2, all the selected variables are significantly crucial at the 95% confidence level. The odds ratios for all the significant variables are also calculated. The odds ratio represents the change in the probability of an explanatory variable influencing the difference in the response variable. For example, the odds of driver yielding is $exp(-0.2354)=0.79$ times lower with each unit (MPH) increase in vehicle speed. Specifically, each unit increase in vehicle speed is associated with a 21% $(1-0.79=0.21)$ reduction in the relative probability of driver yielding, assuming other variables remain fixed (refer to equation (1)). Also, this result exhibits that the odds ratio for the presence of restaurants/bars is 6.24 relative to not present, and 0.197 for the presence of a parking lot than not. Thus, a site with restaurants/bars has 544% $(6.24-1=5.24)$ greater relative driver-yielding probability than a site that does not have restaurants/bars. This result indicates the significance of adjacent land use context on the driver-yielding behavior.
Additionally, longer distances to the nearest park or school are found to negatively impact driver yielding, which has reductions in the relative probability of 98% and 36% in driver yielding rate with each mile increase. Conversely, an increase in the major road’s crossing width positively influences driver yielding, with each additional foot in width leading to a 10.3% increase in the probability.
Intriguingly, the study revealed that the yielding of drivers in the opposite direction adversely affects driver yielding, reducing the relative probability by 80.7%, which contradicts the findings of some prior research [15, 19]. In our analysis, certain factors such as pedestrian volume and weather were found to be insignificant. It is also worth noting that large standard errors for some independent variables might be a consequence of imbalanced data in this study. Specifically, data were collected from 18 locations, but the total number of events varies among these locations, as illustrated in Table 3. This variation may have contributed to the observed disparities and should be considered in interpreting the results.
@@
Fig. 7 visually depicts how the model predicts the probability of driver non-yielding based on one of the independent variables, such as vehicle speed. Vehicle speed emerged as the most crucial variable. As illustrated in Fig. 6(a), the faster the vehicle’s speed at a point near the intersection, the less likely the driver is to yield. This correlation is supported by previous studies [16, 15, 17, 35, 18], and emphasizes the significant role of vehicle speed in forecasting driver behavior. The vehicle speed displays substantial variability between interaction events compared to other predictor variables. It is essential to note that since the vehicle speed is measured at a point near each intersection, causal inference indicates that drivers traveling at higher speeds are less likely to yield.
@@ 4.3 Model Evaluation
The logistic model’s confusion matrix is shown in Table 4.3. It’s evaluated using three metrics: accuracy, precision, and recall, defined in equations (4) and (5). Precision quantifies the ratio of correctly predicted positive instances to the total number of positive predictions made by the model. Recall, on the other hand, is the proportion of correctly predicted positive instances to the total actual positive instances. Notably, the models tend to have more false positives (Type \@slowromancapi@ Error) than false negatives (Type \@slowromancapii@ Error), resulting in higher recall values than precision. This trend indicates that the model is more prone to falsely predict driver yielding when the driver is not actually yielding.
Fig. 8 displays the Receiver Operating Characteristic (ROC) curves, which are plotted with respect to true positive and false positive rates. The ROC curve is a performance measurement for binary classification at various threshold settings. For our model, the area under the ROC curve is 0.88, closely following the left-hand border and top border of the ROC space. This signifies that the model exhibits good performance, adeptly balancing sensitivity and specificity, and also underscores the importance of the selected variables in predicting the outcome of driver-pedestrian interactions.
@@
Lastly, we compare the accuracy in the test data of our logistic regression model with other machine learning models, including Support Vector Machine (0.88), Random Forest (0.885), and Neural Network (0.89). Though the performances are similar, the logistic regression model stands out for its superior interpretability. Due to the sample size constraints, other machine learning models did not surpass logistic regression, leading us to choose it as our primary model.
5 Discussion and Conclusions
In summary, many factors influence whether or not a driver will yield to a pedestrian at an unsignaled intersection. This research endeavors to gather field data from various intersections and carry out exploratory analyses to pinpoint the most crucial factors governing a driver’s yielding behavior. Consequently, we employ the custom-built TIM sensor platform, developed by the Minnesota Traffic Observatory, to obtain vehicle-pedestrian video data. We then construct machine learning models for driver yielding, incorporating distinct features for each model, and utilize these models to evaluate prediction accuracy and identify the most influential variables in determining a driver’s decision to yield to pedestrians. The primary findings of our study can be summarized as follows:
@@ 5 Discussion and Conclusions
  • The construction of the predictive model leveraging selected variables for optimal explanation and potential improvement. These models help in evaluating the variables that most significantly influence driver yielding behavior.
Table 4.2 delineates the specific variables considered in predicting driver yielding. Regression models identify factors such as wider major road crossing width and the presence of restaurants/bars as positive influences, while variables like higher vehicle speed and proximity to parks, schools, parking lots, and opposite-direction vehicles negatively affect yielding
Based on the study’s findings, several actionable recommendations can be made to positively influence driver-yielding behavior. Road width should be minimized while safely accommodating traffic volume, as each additional lane decreases driver yielding by about 35%. Reduced roadway widths may increase driver yielding, provided they maintain sufficient capacity. Pedestrian crossing markings, including continental and standard types, correlate with higher yielding rates and should be included in crossing designs whenever possible. Signs indicating the presence of a crossing should be implemented, as they more than double the probability of driver yielding. Though intersections with bike lanes typically exhibit lower yielding rates, bike lanes’ inherent increase in intersection width might contribute to this finding; thus, additional data is needed to conclusively understand their impact. Consideration of adjacent land use can also help in designing intersections. Intersections in areas with high restaurant density are correlated with higher yielding rates, while those near multi-family housing tend to lower yielding rates. These findings are valuable, though they do not imply causation. These recommendations align with Vision Zero’s goals, fostering a safe and forgiving road environment. Communities across Minnesota and the wider United States can benefit from this research, enhancing ongoing traffic safety initiatives.
Although the dataset covers 18 intersections in Minnesota, its applicability to other regions might differ. Limitations include imbalances in data among locations and a limited scope of studied variables and conditions. Additionally, the absence of certain pedestrian features, such as gender and race, might influence the results.
Future research should consider geographical expansion, additional variables, and diverse machine learning models. Collaborations with urban planners and the integration of other data sources could offer a more comprehensive perspective. Exploration of intervention strategies based on the findings and an analysis of crash frequency in relation to the yielding rate could be valuable next steps.
In conclusion, this study has created and published extensive driver-pedestrian yielding data from 18 Minnesota intersections. As one of the most exhaustive datasets of its kind in the US, it provides significant insights into driver-pedestrian interactions. This research complements ongoing initiatives like Vision Zero and holds broader applicability to regions implementing similar traffic safety strategies. It offers a blueprint for making informed decisions in intersection design, contributing to Vision Zero’s goal of creating a safe road environment. By weaving these findings into traffic planning and design, stakeholders can align with Vision Zero’s overarching aim of eradicating traffic fatalities and severe injuries, embracing a data-driven, multi-disciplinary, and proactive approach.

### v2

@@ 3.3 Data Description
Fig. 3(a) demonstrates the influence of vehicle type on yielding rates. The results indicate that Sports Utility vehicles (SUVs) and sedans are less likely to yield compared to other vehicle types, whereas buses and trucks exhibit the highest yielding rates. We classified pedestrians into six distinct categories to examine differences in driver interactions, as illustrated in Fig. 3(b): A: pedestrian, B: person on a vehicle such as a scooter, C: mix of pedestrian types: scenarios involving more than one pedestrian, D: others, E: person with a dog, and F: person with a stroller or child. The most common type is Category A and the second most common one is pedestrian with a dog (Category E). As exhibited by Fig. 3(c), the drivers are more likely to yield to a group of pedestrians and less likely to yield to a pedestrian alone. In this study, interactions between cyclists and vehicles have been excluded, though the interactions are kept in the published dataset for completion. Fig. 3(d) illustrates the variations in yielding behavior (yielding rate) at different hours of the day. Our findings indicate that drivers are more likely to yield during noon compared to other times throughout the day.
@@
We collected variables that related to the site location to reflect location characteristics so that the results can be used to improve pedestrian walkability. For example, as illustrated in Fig. 4(a), we observed a correlation between the number of bus stops near an intersection and driver yielding rates, for the intersection with bus stops, the more bus stops, the more drivers yield to pedestrian. Notably, intersections with a higher number of bus stops exhibited increased rates of driver yielding. This finding suggests that the presence of bus stops may influence driver awareness and behavior toward pedestrians. Interestingly, this correlation was not observed at intersections without nearby bus stops, indicating a unique relationship between bus stop density and yielding behavior. Regarding the number of lanes on the major road (Fig. 4(b)), drivers are less likely to yield when there are more lanes. Fig. 4(c) reveals a correlation between lower driver-yielding rates and the presence of trees obstructing the drivers’ view of waiting pedestrians. Specifically, the results indicate that the highest driver-yielding rates occur at intersections without any trees blocking the view of crosswalk, while the lowest yielding rate is found at intersections where tree cover is present at most of the intersection corners.
@@
Markings play a crucial role in increasing driver yielding, as demonstrated in Fig. 5(a). Intersections with crosswalk markings exhibit a higher percentage of driver yielding compared to driver non-yielding. Interestingly, sites with standard markings (two parallel lines) had higher yielding rates than those with continental markings, contrary to our expectations. Among the sites, 12 had no markings, 3 had standard markings, and 3 had continental markings. It is important to note the small sample size of sites with standard and continental markings. The presence of a bike lane correlated with a decrease in driver yielding, although only 9.1% of the interactions in the dataset occurred at sites with bike lanes, as shown in Fig. 5(b). As expected, sites with crosswalk signs had higher driver-yielding rates, as depicted in Fig. 5(c).
(a) Comparison of Driver Yielding Rates by Crosswalk Marking Types. ’C’ represents continental markings, ’S’ denotes standard markings with two parallel lines, and ’U’ indicates unmarked crosswalks.
(b) Impact of Bike Lane Presence on Driver Yielding. ’0’ denotes the absence of bike lanes, while ’1’ indicates their presence.
(c) Relationship between crosswalk sign presence and driver yielding. ’0’ represents no crosswalk signs, and ’1’ signifies the presence of crosswalk signs.
@@
Lastly, various adjacent land use types were documented during data extraction, such as distance to a park or school and the presence of multi-family housing. Fig. 6 illustrates the impact of two of these land uses—restaurants/bars and parking lots within 1 block of the intersection. Sites featuring restaurants/bars and/or parking lots constituted a significant portion of the dataset, and these land uses were associated with a higher rate of driver yielding. However, in the case of parking lots, there are still more non-yielding events.
(a) Presence of parking lots vs. yielding. ’0’ denotes the absence of parking lots, while ’1’ indicates their presence.
(b) Presence of restaurants/bars vs. yielding. ’0’ denotes the absence of restaurants/bars, while ’1’ indicates their presence.
@@
Sites Total interactions Close calls Ped Stop midway Yielding rate (%) Total num of ped. Average waiting time (s) Average Traversing time (s) Max waiting time (s) Min waiting time (s)
@@ 4 Modeling Driver Yielding
This section outlines the statistical analysis conducted to model driver-yielding rates at 18 unsignalized intersections under study. Through the application of the developed model, along with comparative analysis against other models, we can identify significant factors influencing driver yielding at these sites and estimate the likelihood of a driver yielding under the prevailing conditions at similar locations.
@@ 4.1 Logistic Regression
Logistic regression is a robust technique for binary classification that yields a model with good interpretability [43, 44]. It models the probability of a binary dependent variable $Y$, which in this case is the driver-yielding response that has a binary outcome - either the driver yields to the pedestrian or fails to yield. We assume a linear relationship between the independent variable $X$ and the log-odds of the event. This linear relationship is shown in (1), where $L$ is the log-odds, $\beta_{i}$ are the model parameters, and $P(Y=1)$ for driver yielding and $1-P(Y=0)$ for driver not yielding are the probabilities of the aforementioned binary outcomes. The logistic regression model takes the form of Equation (2) to estimate the probability of the driver yielding $P(Y=1)$, and conversely, Equation (3) estimates the probability of the driver not yielding $1-P(Y=1)$. For the driver-pedestrian interactions considered, the outcome is either driver-yielding ($Y=1$) or non-yielding ($Y=0$).
@@
where $\beta_{0}$ is the constant, $\beta_{n}$ is the coefficient of the explanatory variable, $x_{n}$ is the predictor variable, and $n$ is the number of features.
The collected interaction data were divided into training and testing data sets, with 80% of the data used for training the model and the remaining 20% reserved as hold-out test data. The parameters used in the logistic model were selected through the stepwise logistic method, which is a technique used to identify the critical explanatory variables. The stepwise process combines forward and backward elimination to select the crucial variables with p-value smaller than $0.05$ [45]. The selection criterion for these predictor variables relies on the Akaike Information Criterion (AIC) [46], a measure that gauges both prediction error and model quality. Recognizing that some variables in this dataset are categorical, we implemented feature engineering to convert these into dummy variables, adopting values of either 0 or 1.
@@ 4.2 Model for Driver-pedestrian Interaction
Table LABEL:tab:log_results lists the variables identified as significant and thus included in the logistic regression model of driver yielding. Based on the variable selection criterion introduced above, the vehicle speed, the yielding action of drivers in the opposite direction, the width of major crossing road, the presence of restaurants/bars/parking lots, the distance to the nearest park, and the distance to the nearest school are all significant. Note that the presence of parking refers to the large street-facing parking lots (not on-street parking) within a 1-block radius of the intersection. Therefore, the logistic regression model presented in Table LABEL:tab:log_results is constructed using the above variables. The coefficients $\beta_{n}$ within the model are determined using maximum-likelihood estimation, an iterative procedure aimed at minimizing the error in the predicted probabilities. These coefficients $\beta_{n}$ articulate the magnitude of the influence each predictor variable has on the binary outcome of driver yielding or not yielding. Additionally, the standard error provides an estimate of each coefficient’s variability, while the Z-score conveys the deviation from the mean in units of standard deviation. A test statistic, the p-value, reflects the probability of obtaining extreme results from the statistical analysis. Specifically, a low p-value signifies a predictor’s likely relevance to the model.
@@
Variables coefficient std err z P>|>|> |z|||| Effect
@@
Based on the result in Table LABEL:tab:log_results, all the selected variables are significantly crucial at the 95% confidence level. The odds ratios for all the significant variables are also calculated. The odds ratio represents the change in the probability of an explanatory variable influencing the difference in the response variable. The odds of driver yielding is $exp(-0.2354)=0.79$ times lower with each unit (MPH) increase in vehicle speed. Specifically, each unit increase in vehicle speed is associated with a 21% $(1-0.79=0.21)$ reduction in the relative probability of driver yielding, assuming other variables remain fixed (refer to equation (1)). Vehicle speed emerged as the most crucial variable based on the Z-score. Therefore, we develop Fig. 7 to visually depict how the model predicts the probability of driver non-yielding based on one of the independent variables: vehicle speed.
In addition, the study revealed that the yielding action of drivers in the opposite direction adversely affects driver yielding, reducing the relative probability by 80.7%. Conversely, wider major road crossings positively influence driver yielding. Each additional foot in road width correlates with a 10.3% increase in the likelihood of drivers yielding.
In the case of built environment-related factors, this result exhibits that the odds ratio for the presence of restaurants/bars is 6.24 relative to not present, and 0.197 for the presence of a parking lot than not. Thus, a site with restaurants/bars has 544% $(6.24-1=5.24)$ greater relative driver-yielding probability than a site that does not have restaurants/bars. Additionally, the proximity of parks or schools appears to have a positive effect on driver-yielding behavior. Specifically, for each mile closer to the nearest park or school, the probability of driver yielding increases by 98% and 36%, respectively. This result indicates the significance of adjacent built environment context on the driver-yielding behavior.
In our analysis, certain factors such as pedestrian group size and weather were found to be insignificant. It is also worth noting that large standard errors for some independent variables might be a consequence of imbalanced data in this study. Specifically, data were collected from 18 locations, but the total number of events varies among these locations, as illustrated in Table 2. This variation may have contributed to the observed disparities and should be considered in interpreting the results.
@@ 4.3 Model Evaluation
The logistic model’s confusion matrix is shown in Table LABEL:tab:confusion_matrix. It’s evaluated using three metrics: accuracy, precision, and recall, defined in equations (4) and (5). Precision quantifies the ratio of correctly predicted positive instances to the total number of positive predictions made by the model. Recall, on the other hand, is the proportion of correctly predicted positive instances to the total actual positive instances. Notably, the models tend to have more false positives (Type \@slowromancapi@ Error) than false negatives (Type \@slowromancapii@ Error), resulting in higher recall values than precision. This trend indicates that the model is more prone to falsely predict driver yielding when the driver is not actually yielding.
@@
Fig. 8 displays the Receiver Operating Characteristic (ROC) curves, which are plotted with respect to true positive and false positive rates. The ROC curve is a performance measurement for binary classification at various threshold settings. For our model, the area under the ROC curve is 0.88 for testing data, closely following the left-hand border and top border of the ROC space. This signifies that the model exhibits good performance, adeptly balancing sensitivity and specificity, and also underscores the importance of the selected variables in predicting the outcome of driver-pedestrian interactions.
@@
Lastly, we compare the accuracy in the test data of our logistic regression model with other machine learning models, including Support Vector Machine (0.88), Random Forest (0.885), and Neural Network (0.89). Though the performances are similar, the logistic regression model stands out for its superior interpretability. Therefore, we choose logistic regression as our primary model.
5 Discussion and Conclusion
In summary, many factors influence whether or not a driver will yield to a pedestrian at an unsignalized intersection. This research endeavors to gather field data from various intersections and carry out in-depth analyses to pinpoint the most crucial factors, including built environment factors, governing a driver’s yielding behavior. Consequently, we employ the custom-built TIM sensor platform, developed by the Minnesota Traffic Observatory, to obtain vehicle-pedestrian video data. We then construct logistic regression models for modeling driver yielding behavior and utilize these models to evaluate prediction accuracy and identify the most influential variables in determining a driver’s decision to yield to pedestrians. The primary findings of our study can be summarized as follows:
@@ 5 Discussion and Conclusion
  • The construction of the predictive model leveraging selected variables for optimal explanation and potential improvement. These models help in evaluating the variables that most significantly influence driver-yielding behavior.
  • Surrounding built environment factors have an influence on driver-yielding behavior, which represents a significant new knowledge contribution to the literature on driver-yielding behavior.
Table LABEL:tab:log_results delineates the specific variables considered in predicting driver yielding. The findings of this study highlight that vehicle speed is the most influential predictor of driver-yielding behavior. The faster the vehicle’s speed at a point near the intersection, the less likely the driver is to yield. This result is intuitive, as high-speed vehicles find it more challenging to stop or slow down for pedestrians compared to lower-speed vehicles, resulting in comparatively lower yielding rates. This result is also supported by previous studies [29, 17, 26, 47, 30], and emphasizes the significant role of vehicle speed in forecasting driver behavior.
The model’s findings indicate that a wider major road crossing width is associated with higher driver-yielding behavior, which differs from the previous study [12]. On the other hand, the study found that when drivers in the opposite direction yield, it reduces the probability of driver yielding, a result contrary to the findings of certain previous studies [17, 18]. One possible explanation for this counterintuitive result is that drivers may assume the pedestrian has already been accommodated by the opposite vehicle and is no longer in their path. This assumption could lead them to believe that further yielding is unnecessary, prompting them to proceed without stopping.
We have categorized the variables into two groups: those related to specific events (e.g., vehicle speed, pedestrian type, interaction type), which are generally beyond immediate control, and those related to the intersection’s built environment. While built environment factors may not be easily modified, our study provides valuable insights that can guide local agencies in making informed decisions for future intersection designs or modifications. It is important to note that at the beginning of our study, we included land use as a feature in our model. However, due to the limited variability in land use characteristics among the 18 intersections we studied, this variable did not prove to be significant.
Built environment factors significantly influence driver-yielding behavior. The presence of restaurants/bars, closeness to schools, and proximity to parks increase driver-yielding rates. These locations typically experience higher levels of pedestrian activity. Therefore, various crash reduction countermeasures, such as speed control, provision of sidewalks, signage, and pedestrian-friendly infrastructure, might already be available in these highly active areas to enhance pedestrian safety, potentially increasing the probability of drivers yielding. However, the presence of parking lots reduces the probability of driver yielding. This is likely due to parking lots reducing the visibility of pedestrians, making it difficult for drivers to anticipate pedestrian arrivals for crossing. Consequently, drivers may not be mentally prepared to slow down or stop for pedestrians, thereby decreasing driver yielding rates.
In case of limitations and further study scope, the dataset covers 18 intersections in Minnesota, its applicability to other regions might vary. Other limitations include data imbalances among locations, a limited scope of studied variables and conditions, and the absence of certain pedestrian features, such as gender and race, which could affect the generalizability of the results. Future research should consider geographical expansion, the incorporation of additional relevant variables, and the use of diverse machine learning models. Data fusion techniques that incorporate multi-source data could be particularly useful in developing models with improved interpretability. Additionally, analyzing crash frequency and severity connecting with the yielding rate, could be valuable next steps.
6 Practical Applications
Based on the study’s findings and discussion, several actionable recommendations can be made to positively influence driver-yielding behavior. Implementation of posted speed limits, installation of speed control measures, and enforcement of speed limits are necessary to control vehicle speed at unsignalized intersections. Replicating effective features found near high pedestrian activity areas (school/park/bar/restaurants), such as pedestrian crossing markings, high-quality pedestrian infrastructure, traffic calming measures, and improved signage and marking, in other pedestrian crash-prone areas, including those near parking lots, can help increase driver yielding behavior. Furthermore, installing pedestrian signals (push buttons) at unsignalized intersections, especially those near parking lots, can increase driver-yielding rates. Furthermore, incorporating comprehensive pedestrian safety training into driver education programs and providing ongoing training for licensed drivers to reinforce the importance of yielding to pedestrians would further enhance driver yielding rates.
These recommendations align with Vision Zero’s goals, fostering a safe and forgiving road environment. The implications of this study extend beyond Minnesota, potentially benefiting traffic safety initiatives throughout the United States. Furthermore, the findings are likely to be valuable in the design and development of automated vehicles in the future.